What is the pH of a 0.0035M KOH solution?

A)2.46
B)5.65
C)8.35
D)11.54
E)None of these

*the answer is D)11.54 but I don't know how to get it.
Answer #1

[OH-]=0.0035 M

pOH =-log[OH] =2.455

pH +pOH = 14

pH=11.545
Answer #2

KOH(aq) -------> K+(aq) + (OH-)(aq)

Now, Sicne KOH is a strong base, therefore it dissociates completely

Hence [OH-] = 0.0035 M

Now, [H+]*[OH-] = 1*10-14

or, [H+] = 2.857*10-12

Thus, pH = -log[H+] = 11.544
Answer #3

[OH-1] = .0035 M

pH = 14 - log (OH-1 )
pH = 14 - log (.0035) = 11.54

D)11.54
Answer #4

[OH-] = .0035 M
pOH = - log (.0035) = 2.455
pH = 14.00 - pOH = 14.00 - 2.455 = 11.54